# CERN-TH-2024-154

### v1

@@ I Introduction
Figure 1: Cyan: excluded region from stellar cooling plus perturbative unitarity of dark Higgs self-coupling, derived in Section II. Magenta: region excluded by overproduction of $A^{\prime}$ from inflation, derived in Section III. Red: Excluded region due to overproduction of $A^{\prime}$ via freeze-in, derived in Section III. Grey: Combination of astrophysical, cosmological and laboratory constraints on dark photons, adapted from Ref. Caputo:2021eaa .
@@ II Higgsed Dark photon constraints
However, there is a theoretical upper limit on $m_{h}^{2}=2\lambda v^{2}$, when written in terms of the VEV and self-coupling $\lambda$ in the potential $V=\lambda(|H|^{2}-v^{2}/2)^{2}$, where $H$ represents the complex field. Perturbative unitarity of $\lambda$ implies that $\lambda<8\pi/3$, while $v$ is related to the dark photon mass by $m_{A^{\prime}}=gv/\sqrt{2}$. These can be tied together by making a generic assumption about the size of the kinetic mixing, in the absence of any special cancellations: the typical contribution from a loop of a heavy particle carrying both U(1) charges gives
@@
$\epsilon<\sqrt{2\over 3}{e\,m_{A^{\prime}}\over 4\pi^{3/2}\,T}$ (2)
@@ III Inflationary generation of $A^{\prime}$ via dark Higgs
Ref. Dror:2018pdh showed how parametric resonance of an oscillating dark Higgs $h$ could efficiently produce $A^{\prime}$ dark matter. Their analysis relied upon arbitrary starting values of $h$ away from its minimum. We note that a definite initial value can be predicted from the stochastic misalignment of $h$ that gets generated by its fluctuations during inflation. Furthermore we employ the simpler mechanism of perturbative decays $h\to A^{\prime}A^{\prime}$ to produce the dark matter. We find that this works in the regime where kinetic mixing is sufficiently small so that $A^{\prime}$ is long-lived compared to the age of the Universe.
During inflation, $h$ acquires a nonzero value through its fluctuations, with amplitude $\langle h^{2}\rangle\sim 3H_{I}^{4}/(8\pi^{2}m_{h}^{2})$ Graham:2018jyp ; Chatrchyan:2022dpy , where where $H_{I}$ is the Hubble rate during inflation, $H_{I}^{2}=\Lambda_{I}^{4}/(3M_{p}^{2})$, with $\Lambda_{I}$ being the energy scale. This corresponds to an energy density
@@
If reheating is efficient, the entropy is initially $s_{i}\sim g_{*}\Lambda_{I}^{3}$. We estimate the abundance of $h$ as
@@
${m_{h}\over m_{A^{\prime}}}\lesssim\sqrt{32\pi\over 3}{e\over 16\pi^{2}\epsilon}\simeq{0.01\over\epsilon}\,.$ (7)
On the other hand, Planck gets an upper limit on the inflation scale, $\Lambda_{I}<1.6\times 10^{16}\,{\rm GeV}$ Planck:2018jri . Hence we cannot saturate the constraint (6) for $\epsilon\lesssim 10^{-11}$, in particular in the flat region of the allowed parameters from Section II. To get the desired $A^{\prime}$ relic density while satisfying (7), we need
$\Lambda_{I}\lesssim 6.3\times 10^{13}\,\epsilon^{-1/5}\,{\rm GeV}\leq 1.6\times 10^{16}\,{\rm GeV}\,.$ (8)
This leads to the constraint $\epsilon\gtrsim 10^{-12}$, which is shown in Fig. 2 (magenta region).
@@ IV Freeze-out by $s$-channel $A^{\prime}$ exchange
by combining Eqs. (10-11) with the working assumption $m_{A^{\prime}}>m_{\chi}$. Since thermal freezeout requires $m_{\chi}\gtrsim 1\,$MeV, it only applies for heavier $A^{\prime}$ than depicted in Fig. 1. In fact, we will show that $m_{A^{\prime}}$ should exceed $\sim 50$ MeV. The excluded region is shown in Fig. 2.
@@
which is much larger than the typical annihilation cross section. Nevertheless, since one generically expects tat $\epsilon\ll g_{\chi}$, it is interesting to determine the minimum ratio $\epsilon/g_{\chi}$ compatible with the combined relic abundance and self-interaction constraints. Combining Eqs.  (10-14), we find that $\epsilon/g_{\chi}$ for thermally produced dark matter is bounded from below by
@@ V Conclusions
Acknowledgements. We thank Miguel Escudero and Gonzalo Alonso-Álvarez for helpful discussions and encouragement. We are grateful to the CERN Theory Department for its hospitality and financial support during our visits. JC is funded by the Natural Sciences and Engineering Research Council (NSERC) of Canada. GH is supported by the the U.S. Department of Energy under award number DE-SC0020250 and DE-SC002026.

### v2

@@ I Introduction
Figure 1: Cyan: excluded region from stellar cooling plus perturbative unitarity of dark Higgs self-coupling, derived in Section II. Magenta: region excluded by overproduction of $A^{\prime}$ from the decays of inflationary misaligned dark Higgs, derived in Section III. Red: Excluded region due to overproduction of $A^{\prime}$ via freeze-in, derived in Section III. Grey: Combination of astrophysical, cosmological and laboratory constraints on dark photons, adapted from Ref. Caputo:2021eaa . Green: Excluded region by cosmic string formation, from Long:2019lwl .
@@ II Higgsed Dark photon constraints
However, there is a theoretical upper limit on $m_{h}^{2}=2\lambda v^{2}$, when written in terms of the VEV and self-coupling $\lambda$ in the potential $V=\lambda(|H|^{2}-v^{2}/2)^{2}$, where $H$ represents the complex field. Perturbative unitarity of $\lambda$ implies that $\lambda<8\pi/3$, while $v$ is related to the dark photon mass by $m_{A^{\prime}}=gv$. These can be tied together by making a generic assumption about the size of the kinetic mixing, in the absence of any special cancellations: the typical contribution from a loop of a heavy particle carrying both U(1) charges gives
@@
$\epsilon<\sqrt{1\over 3}{e\,m_{A^{\prime}}\over 4\pi^{3/2}\,T}$ (2)
@@ III Inflationary generation of $A^{\prime}$ via dark Higgs
Ref. Dror:2018pdh showed how parametric resonance of an oscillating dark Higgs $h$ could efficiently produce $A^{\prime}$ dark matter. Their analysis relied upon arbitrary starting values of $h$ away from its minimum. We note that a definite initial value can be predicted from the stochastic misalignment of $h$ that gets generated by its fluctuations during inflation. Furthermore we employ the simpler mechanism of perturbative decays $h\to A^{\prime}A^{\prime}$ to produce the dark matter. We find that this works in the regime where kinetic mixing is sufficiently small so that $A^{\prime}$ is long-lived compared to the age of the Universe. On the other hand, $h\to A^{\prime}A^{\prime}$ decays are unsuppressed for small $\epsilon$ or $g$; in the limit $g\to 0$, the width goes as $\Gamma\simeq g^{2}m_{h}^{3}/(8\pi m_{A^{\prime}}^{2})=\lambda m_{h}/(4\pi)$ due to the enhanced coupling to longitudinal $A^{\prime}$s Araki:2020wkq . For the parameters of interest, this leads to decays well before big bang nucleosynthesis, except in a small corner (upper left in Fig. 1) that is already excluded by other constraints.
During inflation, $h$ acquires a nonzero value through its fluctuations, with amplitude $\langle h^{2}\rangle\sim 3H_{I}^{4}/(8\pi^{2}m_{h}^{2})$ Graham:2018jyp ; Chatrchyan:2022dpy , where $H_{I}$ is the Hubble rate during inflation, $H_{I}^{2}=\Lambda_{I}^{4}/(3M_{p}^{2})$, with $\Lambda_{I}$ being the energy scale. This corresponds to an energy density
@@
This is only a mean value, and in some Hubble patches, $\langle h^{2}\rangle$ could take a smaller value, in analogy to the misalignment angle for axions accidentally being $\theta\ll 1$. We leave aside this possibility here and simply assume the mean value. If reheating is efficient, the entropy is initially $s_{i}\sim g_{*}\Lambda_{I}^{3}$. We estimate the abundance of $h$ as
@@
${m_{h}\over m_{A^{\prime}}}\lesssim\sqrt{\pi\over 3}{e\over 2\pi^{2}\epsilon}\simeq{0.014\over\epsilon}\,.$ (7)
On the other hand, Planck gets an upper limit on the inflation scale, $\Lambda_{I}<1.6\times 10^{16}\,{\rm GeV}$ Planck:2018jri . Hence we cannot saturate the constraint (6) for $\epsilon\lesssim 10^{-12}$, in particular in the flat region of the allowed parameters from Section II. To get the desired $A^{\prime}$ relic density while satisfying (7), we need
$\Lambda_{I}\lesssim 6.8\times 10^{13}\,\epsilon^{-1/5}\,{\rm GeV}\leq 1.6\times 10^{16}\,{\rm GeV}\,.$ (8)
This leads to the constraint $\epsilon\gtrsim 1.4\times 10^{-12}$, which is shown in Fig. 2 (magenta region).
@@
A potentially efficient alternative mechanism for producing dark photons is through the decay of a network of strings that form during the first-order phase transition when the dark U(1) symmetry breaks spontaneously Redi:2022zkt . For small gauge coupling $g$, the symmetry is approximately global, and decay of $h$ into longitudinal $A^{\prime}$s, or equivalently the Goldstone bosons, is unsuppressed by $g$. Long:2019lwl . This mechanism is efficient in the limit of large string tensions $\mu\cong v^{2}\ln^{2}(m_{h}/m_{A^{\prime}})$, such that
$X\equiv\left(m_{A^{\prime}}\over 10^{-13}\,{\rm eV}\right)^{1/2}{\mu\over(10^{14}\,{\rm GeV})^{2}}\gtrsim 1\,.$ (10)
This is shown as the shaded lower right corner in Fig. 1. For small couplings $\lambda<10^{-3}$ , Ref. Redi:2022zkt showed that isocurvature fluctuations in the produced $A^{\prime}$ particles can lead to strong constraints, but this is opposite to the regime of large $\lambda$ we focus on in the present work.
@@ IV Freeze-out by $s$-channel $A^{\prime}$ exchange
by combining Eqs. (11-12) with the working assumption $m_{A^{\prime}}>m_{\chi}$. Since thermal freezeout requires $m_{\chi}\gtrsim 1\,$MeV, it only applies for heavier $A^{\prime}$ than depicted in Fig. 1. In fact, we will show that $m_{A^{\prime}}$ should exceed $\sim 50$ MeV. The excluded region is shown in Fig. 2.
@@
which is much larger than the typical annihilation cross section. Nevertheless, since one generically expects tat $\epsilon\ll g_{\chi}$, it is interesting to determine the minimum ratio $\epsilon/g_{\chi}$ compatible with the combined relic abundance and self-interaction constraints. Combining Eqs.  (11-15), we find that $\epsilon/g_{\chi}$ for thermally produced dark matter is bounded from below by
@@ V Conclusions
Acknowledgements. We thank Gonzalo Alonso-Álvarez, Andrea Caputo, Miguel Escudero, Andrew Long, Katie Mack, and Michele Redi for helpful discussions. We are grateful to the CERN Theory Department for its hospitality and financial support during our visits. JC is funded by the Natural Sciences and Engineering Research Council (NSERC) of Canada. GH is supported by the the U.S. Department of Energy under award number DE-SC0020250 and DE-SC002026.